OpenStudy (anonymous):

What is the exact value of 5z/√z^5

OpenStudy (lgbasallote):

\[\frac{5z}{\sqrt{z^5}} \implies \frac{5z}{z^2\sqrt z} \implies \frac{5}{z\sqrt z}\] do you still need to rationalize this?

OpenStudy (anonymous):

no this was great help. really struggling with this module. so the final answer should be 5√z/z^2 correct?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thanks!

OpenStudy (lgbasallote):

you're welcome ^_^